# Ground state properties of the Heisenberg-compass model on the square lattice

### v1

@@ Abstract
Compass models provide insights into the properties of Mott-insulating materials that host bond-dependent anisotropic interactions between their pseudospin degrees of freedom. In this article, we explore the classical and quantum ground state properties of one such model relevant to certain layered perovskite materials akin to Ba2IrO4 – namely, the Heisenberg-compass model on the square lattice. We first investigate the ground state phase diagram of this model using classical Monte Carlo simulations. These reveal that the low temperature classical phase diagram is divided into six different classes of long-range ordered phases, including four phases that exhibit an order by disorder selection and two phases that are stabilized energetically. This model admits a special duality transformation, known as the Klein duality, conveniently allowing to map one region of coupling parameters onto another and constraining the phase diagram, and which we exploit in our study. From the analysis of the zero-point energy and the free energy of the spin waves, we find that order by quantum disorder at zero temperature and order by thermal disorder select the same orderings as those found from classical Monte Carlo simulations. We further investigate the quantum ground states of this model using numerical exact diagonalization on small clusters by exploiting the translational symmetry of the square lattice. We obtain a ground state phase diagram bearing close resemblance to that found from the classical analysis.
@@ I Introduction
Lattice models with effective spin-spin interactions offer a fundamental framework for understanding the intriguing behavior of Mott insulating magnets Auerbach (1998); Fazekas (1999). A notable subclass of these models is the family of compass models that have gained significant attention in the context of strongly correlated transition metal (TM) oxides Nussinov and van den Brink (2015). These models are characterized by spatially direction-dependent interactions among their spins. The details of such bond-dependent interactions depend strongly on the symmetries of the underlying system. A well-studied compass model is the Kitaev spin model on the honeycomb lattice Kitaev (2006) which involves three Ising-like spin-spin compass couplings, namely $\cramped{x-x}$, $\cramped{y-y}$, or $\cramped{z-z}$, depending upon the bond orientations of the underlying lattice. This is in contrast to the isotropic Heisenberg interaction or anisotropic interactions that take the same form on every bond in the lattice. Compass interactions are often competing in nature or frustrated, which can lead to a continuous accidental classical ground state degeneracy that is not due to any symmetry of the Hamiltonian. Typically, such degeneracy is not robust to thermal or quantum fluctuations at low energies, and is consequently lifted by these fluctuations, resulting in the stabilization of a magnetically long-range ordered state. This phenomenon is referred to as “order by disorder” Villain et al. (1980); Shender (1982); Henley (1989). Conversely, in certain systems, frustration can prevent ordering down to absolute zero temperature, giving rise to unconventional phases like classical Anderson (1956); Moessner and Chalker (1998) and quantum Canals and Lacroix (1998); Kitaev (2006); Savary and Balents (2016) spin liquids. Compass models thus provide a natural platform for exploring the interplay between conventional notions of order and more exotic disordered states that can arise in frustrated magnets.
@@
Compass models have garnered significant interest due to their applicability across various domains of condensed matter physics. Initially, they were introduced to elucidate a range of physical phenomena in insulating TM oxides with weak spin-orbit coupling, wherein the orbital degrees of freedom of the TM ions couple with each other via compass interactions Kugel´and Khomskiĭ (1982); Feiner et al. (1997); Tokura and Nagaosa (2000); Khomskii and Mostovoy (2003); van den Brink (2004); Mishra et al. (2004). Over time, compass interactions have also been identified in materials with large spin-orbit coupling, where the interactions are not between the orbitals, but rather between pseudospin degrees of freedom Jackeli and Khaliullin (2009). For instance, there has been a surge of efforts in realizing the Kitaev model in real materials with $4d$ and $5d$ TM ions, including $\alpha\textrm{-RuCl}_{3}$ and iridates Jackeli and Khaliullin (2009); Rau et al. (2016); Takagi et al. (2019); Wang et al. (2017). Compass models have also been found relevant in other contexts, such as $p+ip$ superconducting Josephson-junction arrays Xu and Moore (2004); Nussinov and Fradkin (2005), ultracold atoms trapped in optical lattices Wu (2008), settings to safeguard qubits against unwanted decoherence in quantum computing Douçot et al. (2005), higher-form subsystem symmetry breaking Rayhaun and Williamson (2023), dimensional reduction Batista and Nussinov (2005), and strongly interacting topological insulators Tirrito et al. (2022).
@@
Further, exploring material manifestations of compass interactions, individual layers of certain iridium-based perovskites, such as Ba2IrO4 and Tb-substituted Sr2IrO4, offer potential realizations of compass interactions on the square lattice Katukuri et al. (2014); Zhang et al. (2016); Bertinshaw et al. (2019). For example, the Hamiltonian of a single layer in Ba2IrO4 is thought to possess dominant antiferromagnetic Heisenberg and sub-dominant compass exchange Katukuri et al. (2014). Unlike in its cousin Sr2IrO4 Jackeli and Khaliullin (2009), the IrO4 octahedra in Ba2IrO4 do not undergo a staggered rotation Okabe et al. (2011) and the Dzyaloshinskii-Moriya (DM) interaction is forbidden. While Ref. [Katukuri et al., 2014] has provided a detailed analysis of a microscopic model of Ba2IrO4, the complete ground state phase diagram of the in-plane Hamiltonian as a function of Heisenberg and compass interactions has, to the best of our knowledge, not yet been explored. Note that previous theoretical studies have considered more general Heisenberg-compass models on the square lattice Trousselet et al. (2010, 2012), wherein the compass couplings on the $x$ (horizontal) and $y$ (vertical) bonds of the square lattice were not constrained to be identical. These investigations have revealed a rich quantum phase diagram with multiple phase transitions. However, the symmetric compass interaction limit (equal compass coupling on the $x$ and $y$ bonds) in the presence of isotropic Heisenberg exchange, which is relevant to the Ba2IrO4 layers, has not been examined in detail. Thus, investigating the Heisenberg-compass model on the square lattice with symmetric compass interactions is crucial to provide a foundation for understanding the behavior of these perovskites as well as similar spin-orbit coupled magnets on the square lattice that may be synthesized in the future.
@@
In this article, we determine the low-temperature classical and quantum phases of the symmetric compass model on the square lattice in the presence of an additional Heisenberg exchange interaction, hereafter referred to as the “Heisenberg-compass model”. Using classical Monte Carlo simulations and spin-wave analysis, we first establish the classical phase diagram of this model. Our analysis reveals six distinct regimes in the Heisenberg-compass coupling-parameter space, each corresponding to a magnetically ordered phase. These phases are pair-wise related via a unitary transformation – the so-called “Klein duality” familiar from generalized Kitaev models Kimchi and Vishwanath (2014). Of particular interest to the present work are four long-range ordered phases that we identify as arising from order by disorder (ObD) induced by thermal fluctuations, while the remaining two result from conventional energetic selection. Next, we consider the quantum Heisenberg-compass model. We find that in the semi-classical large spin limit, the free energy of quantum spin waves predicts a quantum ObD selection at zero temperature that matches the one found from classical thermal ObD in all four revelant coupling parameter regimes. This persists to finite temperature, with combined quantum and thermal fluctuations preferring the same states at low, but non-zero temperatures. Finally, we tackle the $S=1/2$ limit relevant for real materials. Using exact diagonalization, we determine the quantum ground states at zero temperature, obtaining a qualitatively similar phase diagram to the classical and semi-classical limit, including the selection of the same states as from semi-classical quantum ObD.
@@ II Model
There are a number of important and well-understood limits in the phase diagram of the model. First, at $\xi=0$ and $\xi=\pi$, the Hamiltonian [Eq. (1)] reduces to the well-known Heisenberg antiferromagnet $(J=1)$ and ferromagnet $(J=-1)$, respectively, as marked by ‘HAF’ and ‘HF’ in Fig. 2. At these two special points, the model possesses a global SU(2) symmetry. However, away from these points with a non-zero compass term $(K\neq 0)$, the Hamiltonian no longer has any continuous spin-rotation symmetry. Nevertheless, it still possesses a discrete $C_{4}$ symmetry about the $\bm{\hat{z}}$ axis, which lies normal to the $\bm{\hat{x}}-\bm{\hat{y}}$ plane, and $C_{2}$ symmetries about the $\bm{\hat{x}}$ and $\bm{\hat{y}}$ axes. In the special cases of $\xi=\pi/2$ and $23\pi/$ 2, one has the well-studied antiferromagnetic and ferromagnetic “pure” compass model, respectively Dorier et al. (2005); Mishra et al. (2004); Wenzel and Janke (2008); Nussinov and van den Brink (2015). These two special points are related to one another by the following symmetry; a $\pi$-rotation of the spins about the $\bm{\hat{z}}$ axis on one of the two sublattices of the square lattice maps $K\rightarrow-K$. We therefore refer to both of the ferromagnetic and antiferromagnetic compass models as ‘C’ in Fig. 2. At point C, extra discrete symmetries (special to that compass point) leads to a sub-extensive ground state degeneracy ($\sim 2^{L+1}$ for an $L\times L$ square lattice). In the classical limit at point C, in addition to this symmetry enforced sub-extensive ground state degeneracy, there are accidentally degenerate ground states forming a continuous $O(2)$ manifold Dorier et al. (2005). Interestingly, thermal or quantum fluctuations lift the accidental degeneracy, thus yielding an ObD of colinear states having long-range directional/nematic ordering along the director $\bm{\hat{x}}$ or $\bm{\hat{y}}$ Dorier et al. (2005).
@@ III Klein duality
where $\mathbb{1}$ is the identity rotation and $\mathbb{R}^{\pi}_{\bm{\hat{x}}}$, $\mathbb{R}^{\pi}_{\bm{\hat{y}}}$, and $\mathbb{R}^{\pi}_{\bm{\hat{z}}}$ denote $\pi$-rotations about the $\bm{\hat{x}}$, $\bm{\hat{y}}$, and $\bm{\hat{z}}$ axes, respectively. The rotations, $\Big{\{}\mathbb{1},\mathbb{R}^{\pi}_{\bm{\hat{x}}},\mathbb{R}^{\pi}_{\bm{\hat{y}}},\mathbb{R}^{\pi}_{\bm{\hat{z}}}\Big{\}}$ form an Abelian group, isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, known as the Klein four-group Humphreys (1996). Hence, the transformation in Eq. (2) is referred to as the Klein transformation. Under this transformation, the Hamiltonian [Eq. (1)] maps back to itself, but with modified coupling parameters: $J\rightarrow J^{\prime}\equiv-J$ and $K\rightarrow K^{\prime}\equiv(2J+K)$. Accounting for the change in energy scale, the mapping of the coupling parameters takes the following form,
@@
where the prefactor $1/\sqrt{J^{2}+(2J+K)^{2}}$ takes into account the renormalization of the overall energy scale from the parameter change in $(J,K)\rightarrow(J^{\prime},K^{\prime})$. This transformation exactly maps the properties of the Heisenberg-compass model from one region of the $(J,K)$ parameter space to another one, allowing us to restrict our focus to a smaller subset of the phase diagram and recovering the properties of the model within the remaining regions by an application of the duality.
@@
To visualize this mapping, we connect each pair of points in the parameter space that are dual to one another by dashed lines in Fig. 2. We note from Eq. (3) that when $(2J+K)=0$, the anisotropic Heisenberg-compass model reduces to the isotropic Heisenberg model $(K^{\prime}=0)$. We find that $(2J+K)=0$ admits two solutions: $\xi=\pi-\tan^{-1}(2)$ and $2\pi-\tan^{-1}(2)$. As shown in Fig. 2, $\xi=\pi-\tan^{-1}(2)$ is dual to the pure Heisenberg antiferromagnet (HAF) and labelled as HAF*. Similarly, $\xi=2\pi-\tan^{-1}(2)$ is dual to the pure Heisenberg ferromagnet (HF) and marked as HF*. The Klein duality also reveals that the pure compass points ($\xi=\pi/2,3\pi/2)$ are their self-dual, marked as C*, i.e., they map back to themselves (see Fig. 2).
The entire parameter space is naturally divided into six regimes; Regime-I (HAF – C), Regime-II (C – HAF*), Regime-III (HAF* – HF), Regime-IV (HF – C), Regime-V (C – HF*), and Regime-VI (HF* – HAF). Using the above duality mapping, as depicted in Fig. 2, it is clear that Regime-II, V, and VI are dual to Regime-I, IV, and III, respectively. For simplicity, we shall restrict ourselves only to Regime-I, III, and IV, obtaining the properties of Regime-II, V and VI via the Klein duality.
@@ IV Classical ground states
that flips all the spins on one of the two sublattices of the square lattice. This sublattice spin-flip transformation keeps the classical Hamiltonian [Eq. (1)] invariant if the coupling parameters $(J,K)$ are changed to $(-J,-K)$, i.e., $\xi\rightarrow\xi+\pi$. Similar to the Klein duality [Eq. (2)], this transformation [Eq. (4)] maps different regions of the coupling parameter space onto each other, specifically, relating Region-I, Region-II, and Region-III to Regime-IV, Regime-V, and Regime-VI, respectively. However, it is important to note that reversing the sign of a spin is not a canonical transformation as the Poisson bracket relation for the spin components is not preserved: $\cramped{\{S^{\mu}_{\bm{r}},S^{\nu}_{\bm{r}}\}=\epsilon_{\mu\nu\delta}S^{\delta}_{\bm{r}}}$ changes to $\cramped{\{S^{\mu}_{\bm{r}},S^{\nu}_{\bm{r}}\}=-\epsilon_{\mu\nu\delta}S^{\delta}_{\bm{r}}}$. Consequently, all the properties of the model with parameter set $(J,K)$ are not directly determined by the model with parameter set $(-J,-K)$ resulting from the spin-flip transformation. For instance, the dynamical properties of the transformed Hamiltonian must be different since the equations of motion are not preserved under the sublattice spin-flip transformation [Eq. (4)]. This is in contrast to the canonical Klein duality transformation [Eq. (2)] where all model properties can be mapped exactly between two parameter sets. Nevertheless, as far as the thermodynamic properties of the classical model are concerned, the sublattice spin-flip transformation [Eq. (4)] provides an exact mapping between the parameter sets as the partition function remains unchanged under this transformation. This includes determining the classical ground states and the low-temperature phases of the model, thereby further constraining the classical phase diagram, in addition to the constraints imposed by the Klein duality. Consequently, for the analysis of the classical phase diagram, we may focus on Regime-I and Regime-III, while employing the transformations of Eq. (2) and Eq. (4) to determine the properties of the model in the remaining portions of the phase diagram.

### v2

@@ Abstract
Compass models provide insights into the properties of Mott-insulating materials that host bond-dependent anisotropic interactions between their pseudospin degrees of freedom. In this article, we explore the classical and quantum ground state properties of one such model relevant to certain layered perovskite materials akin to Ba2IrO4 – namely, the Heisenberg-compass model on the square lattice. We first investigate the ground state phase diagram of this model using classical Monte Carlo simulations. These reveal that the low temperature classical phase diagram is divided into six different classes of long-range ordered phases, including four phases that exhibit an order by disorder selection and two phases that are stabilized energetically. This model admits a special duality transformation, known as the Klein duality, conveniently allowing to map one region of coupling parameters onto another and constraining the phase diagram, and which we exploit in our study. From the analysis of the zero-point energy and the free energy of the spin waves, we find that order by quantum disorder at zero temperature and order by thermal disorder select the same orderings as those found from classical Monte Carlo simulations. We further investigate the quantum ground states of this model using numerical exact diagonalization on small clusters by exploiting the translational symmetry of the square lattice. We obtain a ground-state phase diagram bearing close resemblance to that found from the classical analysis.
@@ I Introduction
Lattice models with effective spin-spin interactions offer a fundamental framework for understanding the intriguing behavior of Mott insulating magnets Auerbach (1998); Fazekas (1999). A notable subclass of these models is the family of compass models that have gained significant attention in the context of strongly correlated transition metal (TM) oxides Nussinov and van den Brink (2015). These models are characterized by spatially direction-dependent interactions among their spins. The details of such bond-dependent interactions depend strongly on the symmetries of the underlying system. A well-studied compass model is the Kitaev spin model on the honeycomb lattice Kitaev (2006) which involves three Ising-like spin-spin compass couplings, namely $\cramped{x-x}$, $\cramped{y-y}$, or $\cramped{z-z}$, depending upon the bond orientations of the underlying lattice. This is in contrast to the isotropic Heisenberg interaction or anisotropic interactions that take the same form on every bond in the lattice. Compass interactions are often competing in nature or frustrated, which can lead to a continuous accidental classical ground state degeneracy that is not due to any symmetry of the Hamiltonian. Typically, such degeneracy is not robust to thermal or quantum fluctuations at low energies, and is consequently lifted by these fluctuations, resulting in the stabilization of a magnetically long-range ordered state. This phenomenon is referred to as “order by disorder” Villain et al. (1980); Shender (1982); Henley (1989). Conversely, in certain compass models, frustration can prevent ordering down to absolute zero temperature, giving rise to unconventional phases like spin liquids, for example in Kitaev and generalized Kitaev models Kitaev (2006); Mandal and Surendran (2009); Hermanns and Trebst (2014); O’Brien et al. (2016); Eschmann et al. (2020). Compass models thus provide a natural platform for exploring the interplay between conventional notions of order and more exotic disordered states that can arise in frustrated magnets.
@@
Compass models have garnered significant interest due to their applicability across various domains of condensed matter physics Nussinov and van den Brink (2015). Initially, they were introduced to elucidate a range of physical phenomena in insulating TM oxides with weak spin-orbit coupling, wherein the orbital degrees of freedom of the TM ions couple with each other via compass interactions Kugel´and Khomskiĭ (1982); Feiner et al. (1997); Tokura and Nagaosa (2000); Khomskii and Mostovoy (2003); van den Brink (2004); Mishra et al. (2004). Over time, compass interactions have also been identified in materials with large spin-orbit coupling where the interactions are not between the orbitals, but rather between pseudospin degrees of freedom Jackeli and Khaliullin (2009). For instance, there has been a surge of efforts in realizing the Kitaev model in real materials with $4d$ and $5d$ TM ions, including $\alpha\textrm{-RuCl}_{3}$ and iridates Jackeli and Khaliullin (2009); Rau et al. (2016); Takagi et al. (2019); Wang et al. (2017). Compass models have also been found relevant in other contexts, such as $p+ip$ superconducting Josephson-junction arrays Xu and Moore (2004); Nussinov and Fradkin (2005), ultracold atoms trapped in optical lattices Wu (2008), settings to safeguard qubits against unwanted decoherence in quantum computing Douçot et al. (2005), higher-form subsystem symmetry breaking Rayhaun and Williamson (2023), dimensional reduction Batista and Nussinov (2005), and strongly interacting topological insulators Tirrito et al. (2022).
@@
Further, exploring material manifestations of compass interactions, individual layers of certain iridium-based perovskites, such as Ba2IrO4 and Tb-substituted Sr2IrO4, offer potential realizations of compass interactions on the square lattice Katukuri et al. (2014); Zhang et al. (2016); Bertinshaw et al. (2019). For example, the Hamiltonian of a single layer in Ba2IrO4 is thought to possess dominant antiferromagnetic Heisenberg and sub-dominant compass exchange Katukuri et al. (2014). Unlike in its cousin Sr2IrO4 Jackeli and Khaliullin (2009), the IrO4 octahedra in Ba2IrO4 do not undergo a staggered rotation Okabe et al. (2011) and the Dzyaloshinskii-Moriya (DM) interaction is forbidden. While Ref. [Katukuri et al., 2014] has provided a detailed analysis of a microscopic model of Ba2IrO4, the complete ground state phase diagram of the in-plane Hamiltonian as a function of Heisenberg and compass interactions has, to the best of our knowledge, not yet been explored. Note, however, that previous theoretical studies have considered more general Heisenberg-compass models on the square lattice Trousselet et al. (2010, 2012), wherein the compass couplings on the $x$ (horizontal) and $y$ (vertical) bonds of the square lattice were not constrained to be identical. These investigations have revealed a rich quantum phase diagram with multiple phase transitions. On the other hand, the symmetric compass interaction limit (equal compass coupling on the $x$ and $y$ bonds) in the presence of isotropic Heisenberg exchange, which is relevant to the Ba2IrO4 layers, has not been examined in detail. Thus, investigating the Heisenberg-compass model on the square lattice with symmetric compass interactions is crucial to provide a foundation for understanding the behavior of these perovskites as well as similar spin-orbit coupled magnets on the square lattice that may be synthesized in the future.
@@
In this article, we determine the low-temperature classical and quantum phases of the symmetric compass model on the square lattice in the presence of an additional Heisenberg exchange interaction, hereafter referred to as the “Heisenberg-compass model”. Using classical Monte Carlo simulations and spin-wave analysis, we first establish the classical phase diagram of this model. Our analysis reveals six distinct regimes in the Heisenberg-compass coupling-parameter space, each corresponding to a magnetically ordered phase. These phases are pair-wise related via a unitary transformation – the so-called “Klein duality” familiar from generalized Kitaev models Kimchi and Vishwanath (2014). Of particular interest to the present work are four long-range ordered phases that we identify as arising from order by disorder (ObD) induced by thermal fluctuations, while the remaining two result from conventional energetic selection. Next, we consider the quantum Heisenberg-compass model. We find that in the semi-classical large spin limit, the free energy of quantum spin waves predicts a quantum ObD selection at zero temperature that matches the one found from classical thermal ObD in all four revelant coupling parameter regimes. This persists to finite temperature, with combined quantum and thermal fluctuations preferring the same states at low non-zero temperatures. Finally, we tackle the $S=1/2$ limit relevant for real materials. Using exact diagonalization, we determine the quantum ground states at zero temperature, obtaining a qualitatively similar phase diagram to the classical and semi-classical limit, including the selection of the same states as the ones found from semi-classical quantum ObD.
@@ II Model
There are a number of important and well-understood limits in the phase diagram of the model. First, at $\xi=0$ and $\xi=\pi$, the Hamiltonian [Eq. (1)] reduces to the well-known Heisenberg antiferromagnet $(J=1)$ and ferromagnet $(J=-1)$, respectively, and marked by ‘HAF’ and ‘HF’ in Fig. 2. At these two special points, the model possesses a global SU(2) symmetry. However, away from these points with a non-zero compass term $(K\neq 0)$, the Hamiltonian no longer has any continuous spin-rotation symmetry. Nevertheless, it still possesses a discrete $C_{4}$ symmetry about the $\bm{\hat{z}}$ axis, which lies normal to the $\bm{\hat{x}}-\bm{\hat{y}}$ plane, and $C_{2}$ symmetries about the $\bm{\hat{x}}$ and $\bm{\hat{y}}$ axes. In the special cases of $\xi=\pi/2$ and $23\pi/$ 2, one has the well-studied antiferromagnetic and ferromagnetic “pure” compass model, respectively Dorier et al. (2005); Mishra et al. (2004); Wenzel and Janke (2008); Nussinov and van den Brink (2015). These two special points are related to one another by the following symmetry; a $\pi$-rotation of the spins about the $\bm{\hat{z}}$ axis on one of the two sublattices of the square lattice maps $K\rightarrow-K$. We therefore refer to both of the ferromagnetic and antiferromagnetic compass models as ‘C’ in Fig. 2. At point C, extra discrete symmetries (special to that compass point) leads to a sub-extensive ground state degeneracy ($\sim 2^{L+1}$ for an $L\times L$ square lattice). In the classical limit at point C, in addition to this symmetry enforced sub-extensive ground state degeneracy, there are accidentally degenerate ground states forming a continuous $O(2)$ manifold Dorier et al. (2005). Interestingly, thermal or quantum fluctuations lift this accidental degeneracy, thus yielding an ObD of colinear states having long-range directional/nematic ordering along the director $\bm{\hat{x}}$ or $\bm{\hat{y}}$ Dorier et al. (2005).
@@ III Klein duality
where $\mathbb{1}$ is the identity rotation and $\mathbb{R}^{\pi}_{\bm{\hat{x}}}$, $\mathbb{R}^{\pi}_{\bm{\hat{y}}}$, and $\mathbb{R}^{\pi}_{\bm{\hat{z}}}$ denote $\pi$-rotations about the $\bm{\hat{x}}$, $\bm{\hat{y}}$, and $\bm{\hat{z}}$ axes, respectively. The rotations, $\Big{\{}\mathbb{1},\mathbb{R}^{\pi}_{\bm{\hat{x}}},\mathbb{R}^{\pi}_{\bm{\hat{y}}},\mathbb{R}^{\pi}_{\bm{\hat{z}}}\Big{\}}$ form an Abelian group, isomorphic to $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$, known as the Klein four-group Humphreys (1996). Hence, the transformation in Eq. (2) is referred to as the Klein transformation. Under this transformation, the Hamiltonian [Eq. (1)] maps back to itself, but with modified coupling parameters: $J\rightarrow J^{\prime}\equiv-J$ and $K\rightarrow K^{\prime}\equiv(2J+K)$. The mapping of the coupling parameters takes the following form,
@@
where the prefactor $1/\sqrt{J^{2}+(2J+K)^{2}}$ takes into account the renormalization of the overall energy scale from the parameter change in $(J,K)\rightarrow(J^{\prime},K^{\prime})$. This transformation exactly maps the properties of the Heisenberg-compass model from one region of the $(J,K)$ parameter space to another one, allowing us to restrict our focus to a smaller subset of the phase diagram and recover the properties of the model within the remaining regions by an application of the duality.
@@
To visualize this mapping, we connect each pair of points in the parameter space that are dual to one another by dashed lines in Fig. 2. We note from Eq. (3) that when $(2J+K)=0$, the anisotropic Heisenberg-compass model maps onto the isotropic Heisenberg model $(K^{\prime}=0)$. We find that $(2J+K)=0$ admits two solutions: $\xi=\pi-\tan^{-1}(2)$ and $2\pi-\tan^{-1}(2)$. As shown in Fig. 2, $\xi=\pi-\tan^{-1}(2)$ is dual to the pure Heisenberg antiferromagnet (HAF) and labelled as HAF*. Similarly, $\xi=2\pi-\tan^{-1}(2)$ is dual to the pure Heisenberg ferromagnet (HF) and marked as HF*. The Klein duality also reveals that the pure compass points ($\xi=\pi/2,3\pi/2)$ are their self-dual, marked as C*, i.e., they map back to themselves (see Fig. 2).
The entire parameter space is naturally divided into six regimes; Regime-I (HAF – C), Regime-II (C – HAF*), Regime-III (HAF* – HF), Regime-IV (HF – C), Regime-V (C – HF*), and Regime-VI (HF* – HAF). Using the above duality mapping, as depicted in Fig. 2, it is clear that Regime-II, V, and VI are dual to Regime-I, IV, and III, respectively. For simplicity, we shall restrict ourselves only to Regime-I, III, and IV, obtaining the properties of Regime-II, V and VI by exploiting the Klein duality.
@@ IV Classical ground states
that flips all the spins on one of the two sublattices of the square lattice. This sublattice spin-flip transformation keeps the classical Hamiltonian [Eq. (1)] invariant if the coupling parameters $(J,K)$ are changed to $(-J,-K)$, i.e., $\xi\rightarrow\xi+\pi$. Similar to the Klein duality [Eq. (2)], the transformation in Eq. (4) maps different regions of the coupling parameter space onto each other, specifically, relating Regimes I, II and III to Regimes IV, V, and VI, respectively. However, it is important to note that reversing the sign of a spin is not a canonical transformation as the Poisson bracket relation for the spin components is not preserved: $\cramped{\{S^{\mu}_{\bm{r}},S^{\nu}_{\bm{r}}\}=\epsilon_{\mu\nu\delta}S^{\delta}_{\bm{r}}}$ changes to $\cramped{\{S^{\mu}_{\bm{r}},S^{\nu}_{\bm{r}}\}=-\epsilon_{\mu\nu\delta}S^{\delta}_{\bm{r}}}$. Consequently, not all properties of the model at the spin-flip transformation related parameter sets $(J,K)$ and $(-J,-K)$ are directly related. For instance, the dynamical properties of the transformed Hamiltonian are different from that of the original Hamiltonian since the equations of motion are not preserved under the transformation of Eq. (4). This is in contrast to the canonical Klein duality transformation [Eq. (2)] where all model properties can be mapped exactly between two parameter sets. Nevertheless, as far as the thermodynamic properties of the classical model are concerned, the sublattice spin-flip transformation [Eq. (4)] provides an exact mapping between the parameter sets as the partition function remains unchanged under this transformation. This includes determining the classical ground states and the low-temperature phases of the model, thereby further constraining the classical phase diagram, in addition to the constraints imposed by the Klein duality. Consequently, for the analysis of the classical phase diagram, we may focus on Regime-I and Regime-III, while employing the transformations of Eq. (2) and Eq. (4) to determine the properties of the model in the remaining portions of the phase diagram.